OpenStudy (raheelafzaal):

solve it with sepration. 8cos^2ydx+cosec^2xdy=0

OpenStudy (anonymous):

\[\begin{align*} 8\cos^2y\,dx+\csc^2x\,dy&=0\\\\ \sec^2y\,dy&=-8\sin^2x\,dx \end{align*}\]

OpenStudy (anonymous):

When integrating, you might find this identity useful: \[\sin^2x=\frac{1}{2}-\frac{1}{2}\cos2x\]

OpenStudy (anonymous):

Another way of solving this is by finding an appropriate integrating factor. Notice that multiplying both sides of the ODE by \(\dfrac{\sin^2x}{\cos^2y}\) yields an exact equation, \[8\sin^2x\,dx+\sec^2y\,dy=0\] which gives the same result.